. Find the electric potential of the following configuration at point P. The two short sides of the triangle are the same size = 23 cm.

Q1 = - 5600 nC

Q2 = - 2200 nC


Q1 02
Answer #1


potential due to Q1 , V1 = k*Q1/r1

r1 = sqrt(23^2+23^2) = 32.5 cm = 0.325 m

V1 = -9*10^9*5600*10^-9/0.325

V1 = -155077 V


==============


potential due to Q2 , V2 = k*Q2/r2

r2 = 23 cm = 0.23 m

V2 = -9*10^9*2200*10^-9/0.23

V2 = -86087 V

V = V1 + V2 = -241164 V